OpenStudy (anonymous):

Prove that: [\sum_{n=1}^{\infty}\frac{1}{(2n - 1)^2}] - [\sum_{n=1}^{\infty}\frac{1}{4n^2} = [\sum_{n=1}^{\infty}\frac{1}{n^2}]

sam (.sam.):

\[\huge [\sum_{n=1}^{\infty}\frac{1}{(2n - 1)^2}]-[\sum_{n=1}^{\infty}\frac{1}{4n^2}]=[\sum_{n=1}^{\infty}\frac{1}{n^2}]\]